# Wasserstein Distributionally Robust Shallow Convex Neural Networks

### v1

@@ 5 Physical constraints
where, with respect to sample $j$ and $\forall i\in\llbracket j\rrbracket$, $u_{j}\in\mathbb{R}$ is an upper bound, $l_{j}\in\mathbb{R}$ is a lower bound, $r_{j}\in\mathbb{R}$ is a ramping bound, and (7) is a bounded first-order relation between input and output history with $\underline{\xi}_{j}\in\mathbb{R},\bar{\xi}_{j}\in\mathbb{R},\bm{\eta}_{i}\in\mathbb{R}^{d},\gamma_{i}\in\mathbb{R}$, e.g., a discrete-time dynamical equation. We remark that the WaDiRo-SCNN training problem can be similarly modified to consider physical constraints by adding $\hat{y}_{j}\in\mathcal{Y}_{j}\forall j\in\llbracket N\rrbracket$ to the constraints of (WDR1-SCNNT) and (WDR2-SCNNT).
@@ 6 Numerical study
This section is split into two parts. We first benchmark our approach in a numerical study with data from a controlled environment before showcasing its performance in a real-world power system application.
@@ 6.1.1 Setting
We aim to assess the performance of our model at learning nonlinear functions in heavily corrupted settings, i.e., at being robust against outliers, with a limited amount of training data. To avoid biasing the underlying distribution, we introduce outliers with the sliced Wasserstein distance [41] and a Gaussian noise, in both the training and validation datasets. We choose traditional non-convex benchmark functions: McCormick [42], Picheny-Goldstein-Price (PGP) [43], Keane [44], and Ackley [45]. These functions strike a balance between complexity and tractability. Despite their non-convexity, they exhibit discernable trends and patterns that facilitate learning compared to some other benchmark functions, e.g., Rana or egg-holder [46]. The variation granularity within these functions is moderate ensuring that the learning process is not hindered by excessively fine and extreme fluctuations in the codomain. In a way, this is what we would expect of most real-world regression datasets, e.g., University of California Irvine’s ML repository [47], with the added advantage of being able to control the degree of corruption of the datasets. These functions were generated with the Python library benchmark functions [48] and are illustrated in Figure 1.
@@
We compare our model to an SCNN, a regularized SCNN, a regularized linear regression, a WaDiRo linear regression, and a 4-layer deep $\operatorname{ReLU}$ feedforward neural network (FNN) with batch training and dropout layers. Each of these approaches uses a $\ell_{1}$-loss function. We use simulated annealing [49] for hyperparameter optimization with hyperopt [50]. We run 400 trials per ML model, per benchmark function, except for the deep FNN which we limit to 100 trials to mitigate its substantial training time. We split the data such that the training data accounts for 60% of the whole dataset and we split the rest evenly between the validation and the testing phases. We note that the testing data is uncorrupted to measure the ability of each method to learn the true underlying functions without having access to it. We use the mean average error (MAE) for hyperparameter tuning. We also measure the mean squared error (MSE) for additional performance insights. Our whole experiment setup is available on our GitHub page [^1] https://github.com/jupall/WaDiRo-SCNN for reproducibility and is delineated in Appendix B. In Experiment A, the outliers are introduced solely in the training data while they are split randomly between training and validation for all the other experiments. Table 1 details the experiment contexts.

### v2

@@ 5 Physics-constrained shallow convex neural networks
where in (6), (7), and (8), with respect to sample $j$ and $\forall i\in\llbracket j\rrbracket$, $u_{j}\in\mathbb{R}$ is an upper bound, $l_{j}\in\mathbb{R}$ is a lower bound, and $r_{j}\in\mathbb{R}$ is a ramping bound. Additionally, (8) is a bounded first-order relation between input and output history with $\underline{\xi}_{j}\in\mathbb{R},\bar{\xi}_{j}\in\mathbb{R},\bm{\eta}_{i}\in\mathbb{R}^{d},\gamma_{i}\in\mathbb{R}$, e.g., a discrete-time dynamical equation. These examples are linear or easily linearizable but we stress that any convex constraint can be included. In the case of non-differentiable convex functions, the numerical solver must be chosen accordingly or a subgradient method be implemented Boyd and Park (2014). We remark that the WaDiRo-SCNN training problem can be similarly modified to consider physical constraints by adding $\hat{y}_{j}\in\mathcal{Y}_{j}\ \forall j\in\llbracket N\rrbracket$ to the constraints of (WDR1-SCNNT) and (WDR2-SCNNT).
Proposition 4.
Assume that mappings (2) and (3) are exact, i.e., $m\geq m^{\star}$, then (PCSCNNT) is exactly equivalent to adding convex physical constraints to (SNNT) for any additive convex loss function $\mathcal{L}$, e.g., $\|\cdot\|_{1}$ and $\|\cdot\|_{2}^{2}$.
Proof: To prove Proposition 4 we utilize the exact mappings of the original problem and show that the integration of physical constraints into (SNNT) yields to (PCSCNNT). We start by stating the physics-constrained non-convex problem:
$p^{\star}_{\mathcal{Y}}=$ $\min_{\mathbf{W}_{1},\mathbf{w}_{2}}$ $\sum_{j=1}^{N}\mathcal{L}\left(\sum_{i=1}^{m}\max\left\{\mathbf{x}_{j}^{\top}\mathbf{W}_{1i},0\right\}w_{2i},y_{j}\right)$
s.t. $\sum_{i=1}^{m}\max\left\{\mathbf{x}_{j}^{\top}\mathbf{W}_{1i},0\right\}w_{2i}\in\mathcal{Y}_{j}\quad\forall j\in\llbracket N\rrbracket.$
Substituting mappings (2) and (3), and adding the cone constraints, we have:
$p^{\star}_{\mathcal{Y}}=$ $\min_{\bm{\nu},\bm{\omega}}$ $\sum_{j=1}^{N}\mathcal{L}\left(\sum_{i=1}^{m^{\star}/2}\max\left\{\mathbf{x}_{j}^{\top}\frac{\bm{\nu_{i}}}{\sqrt{\|\bm{\nu_{i}}\|_{2}}},0\right\}\sqrt{\|\bm{\nu_{i}}\|_{2}}+\max\left\{\mathbf{x}_{j}^{\top}\frac{\bm{\omega_{i}}}{\sqrt{\|\bm{\omega_{i}}\|_{2}}},0\right\}(-\sqrt{\|\bm{\omega_{i}}\|_{2}}),y_{j}\right)$
s.t. $\sum_{i=1}^{m^{\star}/2}\max\left\{\mathbf{x}_{j}^{\top}\frac{\bm{\nu_{i}}}{\sqrt{\|\bm{\nu_{i}}\|_{2}}},0\right\}\sqrt{\|\bm{\nu_{i}}\|_{2}}+\max\left\{\mathbf{x}_{j}^{\top}\frac{\bm{\omega_{i}}}{\sqrt{\|\bm{\omega_{i}}\|_{2}}},0\right\}(-\sqrt{\|\bm{\omega_{i}}\|_{2}})\in\mathcal{Y}_{j}\quad\forall j\in\llbracket N\rrbracket$
$(2\text{1}(\mathbf{x}_{j}^{\top}\mathbf{s}_{i}\geq 0)-1)\mathbf{x}_{j}^{\top}\bm{\nu_{i}}\geq 0\quad\forall j\in\llbracket N\rrbracket,\forall i\in\left\llbracket\frac{m^{\star}}{2}\right\rrbracket$
$(2\text{1}(\mathbf{x}_{j}^{\top}\mathbf{s}_{i}\geq 0)-1)\mathbf{x}_{j}^{\top}\bm{\omega_{i}}\geq 0\quad\forall j\in\llbracket N\rrbracket,\forall i\in\left\llbracket\frac{m^{\star}}{2}\right\rrbracket.$
Using the hyperplane arrangement from which we obtained the possible activation patterns, we can show that if there is an activation for a specific $ij$, then $\max\left\{\mathbf{x}_{j}^{\top}\frac{\bm{\nu_{i}}}{\sqrt{\|\bm{\nu_{i}}\|_{2}}},0\right\}\sqrt{\|\bm{\nu_{i}}\|_{2}}=\mathbf{x}_{j}^{\top}\bm{\nu_{i}},\text{ and}\ \max\left\{\mathbf{x}_{j}^{\top}\frac{\bm{\omega_{i}}}{\sqrt{\|\bm{\omega_{i}}\|_{2}}},0\right\}(-\sqrt{\|\bm{\omega_{i}}\|_{2}})=-\mathbf{x}_{j}^{\top}\bm{\omega_{i}}$, and otherwise both equals 0. Thus, we can completely substitute the $\operatorname{ReLU}$ operator by $d_{ji}=\text{1}(\mathbf{x}_{j}^{\top}\mathbf{s}_{i}\geq 0)$, leading to:
$p^{\star}_{\mathcal{Y}}=$ $\min_{\bm{\nu},\bm{\omega}}$ $\sum_{j=1}^{N}\mathcal{L}\left(\sum_{i=1}^{m^{\star}/2}d_{ji}\mathbf{x}_{j}^{\top}(\bm{\nu}_{i}-\bm{\omega}_{i}),y_{j}\right)$
s.t. $\sum_{i=1}^{m^{\star}/2}d_{ji}\mathbf{x}_{j}^{\top}(\bm{\nu}_{i}-\bm{\omega}_{i})\in\mathcal{Y}_{j}\quad\forall j\in\llbracket N\rrbracket$
$(2d_{ji}-1)\mathbf{x}_{j}^{\top}\bm{\nu_{i}}\geq 0\quad\forall j\in\llbracket N\rrbracket,\forall i\in\left\llbracket\frac{m^{\star}}{2}\right\rrbracket$
$(2d_{ji}-1)\mathbf{x}_{j}^{\top}\bm{\omega_{i}}\geq 0\quad\forall j\in\llbracket N\rrbracket,\forall i\in\left\llbracket\frac{m^{\star}}{2}\right\rrbracket,$
from which we find (PCSCNNT). $\square$
We remark that a similar approach can be used for the regularized problem.
6 Post-Training Certification
As mentioned previously, there is an appeal in having ML model architectures that are easily integrable in post-training verification frameworks. We now showcase how to integrate SCNNs in such frameworks by proposing a local-Lipschitz stability certification program to evaluate the stability induced by different SCNN training programs, viz., (WDR1-SCNNT), (WDR2-SCNNT), and (SCNNT). We certify the worst-case output deviation of SCNNs when subject to a small perturbation $\varepsilon\in\mathcal{E}$ in their feature space. This methodology is inspired by Huang et al. (2021) on $\operatorname{ReLU}$ feedforward deep neural networks, but tailored to SCNNs thanks to the decoupled formulation. Recall the definition of the Lipschitz constant for a function $f:\mathcal{X}\to\mathbb{R}$. A function is $K$-Lipschitz if there exists a constant $0<K<+\infty$ such that :
$\frac{|f(\mathbf{x}_{1})-f(\mathbf{x}_{2})|}{|\mathbf{x}_{1}-\mathbf{x}_{2}|}\leq K,$
for any $\mathbf{x}_{1},\mathbf{x}_{2}\in\mathcal{X}$. In our case, we consider only neighboring points by defining $\mathbf{x}_{2}=\{\mathbf{x}_{1}+\bm{\varepsilon}|\bm{\varepsilon}\in\mathcal{E},\mathbf{x}_{1}\in\mathcal{X},\mathcal{E}\text{ is a small convex set}\}$, thus the term local-Lipschitz. We define the $\varepsilon$-stability of a function $f$ by $L_{\varepsilon}$ such that:
$L_{\varepsilon}=\max_{\mathbf{x}_{1},\bm{\varepsilon}:\mathbf{x}_{1}\in\mathcal{X},\bm{\varepsilon}\in\mathcal{E}}|f(\mathbf{x}_{1})-f(\mathbf{x}_{1}+\bm{\varepsilon})|,$
which is visibly linked to the local-Lipschitz bound.
Proposition 5.
Let $\bm{\beta}^{\star}$ be the optimal weights obtained by an SCNN training program and a set of sampling vectors $\mathcal{S}$ obtained before training. The worst possible output deviation $0<L_{\varepsilon}<+\infty$ of the SCNN for any perturbation vector $\bm{\varepsilon}$, from a small convex set $\mathcal{E}$, on any input from the input space can be obtained by a mixed-integer convex program (MICP).
Proof: Let $M\in\mathbb{R}_{>0}$ a large strictly positive scalar, $\mathbf{x}\in\mathcal{X}$ any possible input from feature space $\mathcal{X}\subseteq\mathbb{R}^{d}$, $\mathbf{x}_{\varepsilon}\in\mathcal{X}$ a perturbed input close to $\mathbf{x}$. The post-training verification problem can be formulated as:
$L_{\varepsilon}=$ $\max_{\hat{y},\hat{y}_{\varepsilon},\mathbf{x},\bm{\varepsilon},\mathbf{A},\mathbf{B},\bm{\gamma},\bm{\eta}}$ $|\hat{y}-\hat{y}_{\varepsilon}|$ (9)
[…]
$\hat{\mathbf{x}}$ $=\left(\operatorname{vec}(\mathbf{A}),\operatorname{vec}(\mathbf{A})\right)\quad$ (11)
$\mathbf{x}^{\top}\mathbf{s}_{i}$ $\leq M\gamma_{i}$ $\forall i\in\llbracket P\rrbracket$ (12)
$\mathbf{x}^{\top}\mathbf{s}_{i}$ $\geq-M(1-\gamma_{i})$ $\forall i\in\llbracket P\rrbracket$ (13)
$x_{k}$ $\leq a_{ki}+M(1-\gamma_{i})$ $\forall i\in\llbracket P\rrbracket,\ \forall k\in\llbracket d\rrbracket$ (14)
$x_{k}$ $\geq a_{ki}-M(1-\gamma_{i})$ $\forall i\in\llbracket P\rrbracket,\ \forall k\in\llbracket d\rrbracket$ (15)
$a_{ki}$ $\leq M\gamma_{i}$ $\forall i\in\llbracket P\rrbracket,\ \forall k\in\llbracket d\rrbracket$ (16)
$a_{ki}$ $\geq-M\gamma_{i}$ $\forall i\in\llbracket P\rrbracket,\ \forall k\in\llbracket d\rrbracket$ (17)
$\bm{\gamma}$$\in\{0,1\}^{P}$ (18)
$\hat{y}_{\varepsilon}$ $=\bm{\beta}^{\star\top}\hat{\mathbf{x}}_{\varepsilon}$ (19)
$\hat{\mathbf{x}}_{\varepsilon}$ $=\left(\operatorname{vec}(\mathbf{B}),\operatorname{vec}(\mathbf{B})\right)\quad$ (20)
${\mathbf{x}}_{\varepsilon}$ $=(\mathbf{x}+\bm{\varepsilon})$ (21)
$\mathbf{x}_{\varepsilon}^{\top}\mathbf{s}_{i}$ $\leq M\eta$ $\forall i\in\llbracket P\rrbracket$ (22)
$\mathbf{x}_{\varepsilon}^{\top}\mathbf{s}_{i}$ $\geq-M(1-\eta_{i})$ $\forall i\in\llbracket P\rrbracket$ (23)
$x_{\varepsilon,k}$ $\leq b_{ki}+M(1-\eta_{i})$ $\forall i\in\llbracket P\rrbracket,\ \forall k\in\llbracket d\rrbracket$ (24)
$x_{\varepsilon,k}$ $\geq b_{ki}-M(1-\eta_{i})$ $\forall i\in\llbracket P\rrbracket,\ \forall k\in\llbracket d\rrbracket$ (25)
$b_{ki}$ $\leq M\eta_{i}$ $\forall i\in\llbracket P\rrbracket,\ \forall k\in\llbracket d\rrbracket$ (26)
$b_{ki}$ $\geq-M\eta_{i}$ $\forall i\in\llbracket P\rrbracket,\ \forall k\in\llbracket d\rrbracket$ (27)
$\bm{\eta}$$\in\{0,1\}^{P}$ (28)
$\mathbf{s}_{i}$ $\in\mathcal{S}$ $\forall i\in\llbracket P\rrbracket$ (29)
$\bm{\varepsilon}$ $\in\mathcal{E}$ (30)
$\mathbf{x}$ $\in{\Xi},$ (31)
$\mathbf{x},{\mathbf{x}}_{\varepsilon}$ $\in\mathcal{X}$ (32)
where $\mathbf{A}=(\mathbf{a}_{1},\mathbf{a}_{2},\ldots,\mathbf{a}_{P})$ and $\mathbf{B}=(\mathbf{b}_{1},\mathbf{b}_{2},\ldots,\mathbf{b}_{P})$ are intermediate variables. $\mathcal{E},\Xi\subset\mathbb{R}^{d}$ are convex sets that ensure that $\bm{\varepsilon}$ is finite and small and that $\mathbf{x}$ is finite, respectively, e.g., box constraints on each element or bounded Euclidian norms. Constraints (10)--(18) and (20)--(28) are used to represent the $\operatorname{ReLU}$-like activation patterns of the SCNN for the unperturbed and perturbed output, respectively. Specifically, matrices $\mathbf{A}$ and $\mathbf{B}$ model $\mathbf{xd}^{\top}$ and $\mathbf{x_{\varepsilon}d_{\varepsilon}}^{\top}$ with constraints (14)--(17) and (24)--(27), respectively. Constraints (12), (13), (22), and (23) embed the following definitions: $\mathbf{d}^{\top}=\begin{bmatrix}\text{1}(\mathbf{x}^{\top}\mathbf{s}_{1}\geq 0)&\text{1}(\mathbf{x}^{\top}\mathbf{s}_{2}\geq 0)&\ldots&\text{1}(\mathbf{x}^{\top}\mathbf{s}_{P}\geq 0)\end{bmatrix}$ and $\mathbf{d}_{\varepsilon}^{\top}=\begin{bmatrix}\text{1}(\mathbf{x}_{\varepsilon}^{\top}\mathbf{s}_{1}\geq 0)&\text{1}(\mathbf{x}_{\varepsilon}^{\top}\mathbf{s}_{2}\geq 0)&\ldots&\text{1}(\mathbf{x}_{\varepsilon}^{\top}\mathbf{s}_{P}\geq 0)\end{bmatrix}$. Finally, the non-concave objective function can be linearized using another binary variable:
$\max_{\hat{y},\hat{y}_{\varepsilon},\mathbf{x},\bm{\varepsilon},\mathbf{A},\mathbf{B},\bm{\gamma},\bm{\eta}}\alpha$ (33)
[…]
$\hat{y}_{\varepsilon}-\hat{y}$ $\geq M(\zeta-1)$ (35)
$\alpha$ $\leq\hat{y}_{\varepsilon}-\hat{y}+M(1-\zeta)$ (36)
$\alpha$ $\geq\hat{y}_{\varepsilon}-\hat{y}-M(1-\zeta)$ (37)
$\alpha$ $\leq\hat{y}-\hat{y}_{\varepsilon}+M\zeta$ (38)
$\alpha$ $\geq\hat{y}-\hat{y}_{\varepsilon}-M\zeta$ (39)
$\zeta$$\in\{0,1\}$ (40)
[…]
$\square$
Besides pure certification, this formulation is interesting because it could be used in a pessimistic bilevel program, see Dempe (2002), that integrates SCNN training and certification into the same problem. We leave this opening for future work. We remark that $M$ can be tuned separately for each constraint to be as tight as possible and that recent works have greatly accelerated similar procedures for deep $\operatorname{ReLU}$ neural networks, e.g., see Zhou et al. (2024).
7 Outlier generation on benchmark functions
In this paper, to generate empirical distributions with complex patterns, we propose to use the mathematical expressions of traditional non-convex benchmark functions as a starting point. These functions are normally used to evaluate optimization solvers because of their interesting patterns and their generalizability to $n$ dimensions.
We use a methodology inspired by the work of Pallage et al. (2024) to generate outliers. They propose a method based on the sliced-Wasserstein (SW) distance to remove out-of-distribution outlier points from training sets. In this work, we do the opposite and use a SW-based method to ensure that introduced outliers are out-of-distribution with respect to the original data distribution.
Computing the Wasserstein distance is generally of high computational complexity, but the one-dimensional Wasserstein distance is an exception as it possesses a closed-form solution. The SW distance exploits this special case by first computing, through linear projections, an infinite amount of one-dimensional representations for each original $n$-dimensional distribution. It then computes the average Wasserstein distance between their one-dimensional projections (Kolouri et al. 2019). In practice, the order-$t$ SW distance’s approximation under a Monte Carlo scheme of $L$ samples is defined as:
$SW_{\|\cdot\|,t}\left(\mathbb{U},\mathbb{V}\right)\approx\left(\frac{1}{L}\sum_{l=1}^{L}W_{\|\cdot\|,t}\left(\mathfrak{R}\mathbb{U}\left(\cdot,\theta_{l}\right),\mathfrak{R}\mathbb{V}\left(\cdot,\theta_{l}\right)\right)^{t}\right)^{1/t},$
where $\mathfrak{R}\mathbb{D}(\cdot,\bm{\theta}_{l})$ is a single projection of the Radon transform of distribution function $\mathbb{D}$ over a fixed sample $\bm{\theta}_{l}\in\mathbb{S}^{d-1}=\{\bm{\theta}\in\mathbb{R}^{d}|\theta_{1}^{2}+\theta_{2}^{2}+\cdots+\theta_{d}^{2}=1\}$ $\forall l\in\llbracket L\rrbracket$. Note that $\mathbb{S}^{d-1}$ is often referred to as the $(d-1)$-dimensional unit sphere as it encloses every $d$-dimensional unit-norm vector in a sphere of radius 1. It can be shown that the WD and SWD are closely related (Bonnotte 2013, Theorem 5.1.5). We refer interested readers to Bonneel et al. (2015) and Kolouri et al. (2019) as they offer deeper mathematical insights.
Our procedure is as follows. Consider an empirical distribution of $N$ samples ^$\hat{\mathbb{P}}_{N}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\mathbf{z}_{i}}(\mathbf{z})$ obtained by sampling uniformally a non-convex benchmark function denoted by $\mathbb{P}$. We use the notation ^$\hat{\mathbb{P}}_{N+m}$ to denote the empirical distribution with $m$ added samples from $\mathbb{P}$ and ^$\hat{\mathbb{P}}_{N}\cup\mathcal{O}_{m}$ to represent the union of the empirical distribution with a set of $m$ outlier samples $\mathcal{O}_{m}$. Let $\varphi_{m}>0$ represent a threshold SW distance for sample $m$. The SW distance deals with distributions of equal weights and must account for present outliers when generating a new one. Assuming we want to generate a total of $\overline{m}\in\mathbb{Z}_{>0}$ out-of-sample outliers, the iterative procedure is based on:
^$\hat{\mathbb{P}}_{N}\cup\mathcal{O}_{m}=\left\{\hat{\mathbb{P}}_{N}\cup\mathcal{O}_{m-1}\cup\{\tilde{\mathbf{z}}\}\ \middle|\tilde{\mathbf{z}}\in\mathcal{Z},\left(SW_{\|\cdot\|,t}(\hat{\mathbb{P}}_{N+1}\cup\mathcal{O}_{m-1},\hat{\mathbb{P}}_{N}\cup\mathcal{O}_{m-1}\cup\{\tilde{\mathbf{z}}\})\geq\varphi_{m}\right)\right\}\quad m\in\llbracket\overline{m}\rrbracket,$
where $\mathcal{O}_{0}=\emptyset$ is the empty set. This procedure ensures that each new outlier creates a shift of at least $\varphi$ with respect to the empirical distribution and previous outliers 1, $11,\ 2,\ldots,m-$ 1. We must consider previous outliers to ensure we do not create any new involuntary pattern in the data. Because we do not make a single assumption on $\mathbb{P}$, this methodology is adaptable to generate outliers on any function.
@@ 8 Numerical study
This section is split into two parts. We first benchmark our approach in a numerical study with data from a controlled environment before showcasing its performance in a real-world VPP application.
@@ 8.1.1 Setting
We aim to assess the performance of our model at learning nonlinear functions in heavily corrupted settings, i.e., at being robust against outliers, with a limited amount of training data. To avoid biasing the underlying distribution, we introduce outliers with the SW distance as presented in section 7 and a Gaussian noise, in both the training and validation datasets. We choose traditional non-convex benchmark functions: McCormick (Adorio 2005), Picheny-Goldstein-Price (PGP) (Picheny et al. 2013), Keane (Keane 1994), and Ackley (Ackley 2012). These functions strike a balance between complexity and tractability. Despite their non-convexity, they exhibit discernable trends and patterns that facilitate learning compared to some other benchmark functions, e.g., Rana or egg-holder (Whitley et al. 1996). The variation granularity within these functions is moderate ensuring that the learning process is not hindered by excessively fine and extreme fluctuations in the codomain. In a way, this is what we would expect of most real-world regression datasets, e.g., University of California Irvine’s ML repository (Kelly et al. 2023), with the added advantage of being able to control the degree of corruption of the datasets. These functions were generated with the Python library benchmark functions (Baronti and Castellani 2024) and are illustrated in Figure 2.
@@
We compare our model to an SCNN, a regularized SCNN, a regularized linear regression, a WaDiRo linear regression, and a 4-layer deep $\operatorname{ReLU}$ feedforward neural network (FNN) with batch training and dropout layers. Each of these approaches uses a $\ell_{1}$-loss function. WaDiRo-SCNNs and SCNNs are modelled and optimized with cvxpy (Diamond and Boyd 2016) and with Clarabel (Goulart and Chen 2024) as the solver. We then extract the weights and encode them into pytorch’s neural network module (Paszke et al. 2019). Similarly, the WaDiRo linear regression and the regularized linear regression are optimized with cvxpy and Clarabel, while the deep FNN is fully implemented with pytorch. We use simulated annealing (Ingber 1993) for hyperparameter optimization with hyperopt (Bergstra et al. 2015). We run 400 trials per ML model, per benchmark function, except for the deep FNN which we limit to 100 trials to mitigate its substantial training time. We split the data such that the training data accounts for 60% of the whole dataset and we split the rest evenly between the validation and the testing phases. We note that the testing data is uncorrupted to measure the ability of each method to learn the true underlying functions without having access to it. We use the mean average error (MAE) for hyperparameter tuning. We also measure the root mean squared error (RMSE) for additional performance insights. Our whole experiment setup is available on our GitHub page [^1] https://github.com/jupall/WaDiRo-SCNN for reproducibility and is delineated in Appendix B. In Experiment A, the outliers are introduced solely in the training data while they are split randomly between training and validation for all the other experiments. Table 1 details the experiment contexts.